OpenStudy (anonymous):

use the interactive matrix tool to solve the system of equations. 3x+4y=18 -2x+5y=34

OpenStudy (matheducatormcg):

do they mean on your calculator? or some sort of applet?

OpenStudy (anonymous):

im not sure...it just says that on my homework?

OpenStudy (anonymous):

i think its something like this..... http://answers.yahoo.com/question/index?qid=20100519152741AAliq6f

OpenStudy (matheducatormcg):

you know anything about matrices? (plural of matrix)

OpenStudy (anonymous):

NOPE! lol

OpenStudy (matheducatormcg):

well, i don't know that I can properly teach that concept in this medium. this some online course?

OpenStudy (matheducatormcg):

you got a TI-83 or any higher graphing calculator?

OpenStudy (matheducatormcg):

i could walk you step by step with one of those.

OpenStudy (anonymous):

yes its an online course

OpenStudy (anonymous):

no i dont :x

OpenStudy (matheducatormcg):

if you have to know them, study here http://tutorial.math.lamar.edu/Classes/LinAlg/SystemEqns_Matrices.aspx here are some other sites to help you http://www.coolmath.com/algebra/index.html http://khanacademy.org http://patrickjmt.com good luck in your quest and seriously check these sites.

OpenStudy (matheducatormcg):

i'll solve this one matrix styley! \[\left[\begin{matrix}3 & 4 \\ -2 & 5\end{matrix}\right]\left[\begin{matrix}x \\ y\end{matrix}\right]=\left[\begin{matrix}18 \\ 34\end{matrix}\right]\] Let A be the 2x2 matrix and let C be the 2x1 matrix it is equal to. By matrices we can find what x and y equal to make this true. \[AX=C\] \[A^{-1}AX=A^{-1}C\] \[X=A^{-1}C\] which will be a 2x1 matrix. if \[\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\] is your matrix, then \[A^{-1}=\frac{1}{det(A)}\left[\begin{matrix}d & -b \\ -c & a\end{matrix}\right]\] Where \[det(A)=ad-bc\] so, \[A^{-1}=\frac{1}{23}\left[\begin{matrix}5 & -4 \\ 2 & 3\end{matrix}\right]\] \[A^{-1}=\left[\begin{matrix}\frac{5}{23} & \frac{-4}{23} \\ \frac{2}{23} & \frac{3}{23} \end{matrix}\right]\] Now let's find out what \[A^{-1}C\] is equal to. \[\left[\begin{matrix}\frac{5}{23} & \frac{-4}{23} \\ \frac{2}{23} & \frac{3}{23} \end{matrix}\right]\left[\begin{matrix}18 \\ 34\end{matrix}\right]\] \[\left[\begin{matrix}-2 \\ 6\end{matrix}\right]\] therefore, x=-2, y=6. your ordered pair is (-2,6)

OpenStudy (matheducatormcg):

fun.

OpenStudy (anonymous):

CONFUSING